OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. These are the options : a. y = -3x - 13 b. y = -3x + 11 c. y = -3x + 13 d. y = -3x + 1

jigglypuff314 (jigglypuff314):

-2 = (3•3) + b solve for b

OpenStudy (anonymous):

b = 9

OpenStudy (anonymous):

I think

jigglypuff314 (jigglypuff314):

not quite :) and are you sure there aren't any typos in your options or question?

OpenStudy (anonymous):

My bad these are the options lol a. y = 3x + 7 b. y = 3x - 11 c. y = -3x + 11 d. y = -3x - 12

jigglypuff314 (jigglypuff314):

cool :) then try -2 = (3•3) + b -2 = 9 + b solve for b

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

for b

jigglypuff314 (jigglypuff314):

mmm close! :) -2 = 9 + b subtract 9 from both sides -9 -9 b = -2 - 9 = ? what is -2 - 9 = ?

OpenStudy (anonymous):

-2 -9 = -11

jigglypuff314 (jigglypuff314):

yep :) so we got that b = -11 so y = 3x - 11

OpenStudy (anonymous):

Ok thanks :] so i'm trying to learn how to still do it i get confused .. I got another question ima try doing it by myself if i cant get it i'll post it here

jigglypuff314 (jigglypuff314):

okay :) good luck!